### v1

# Extending the European Competence Framework for Quantum Technologies: new proficiency triangle and qualification profiles

Abstract
With the increasing industrial relevance of quantum technologies (QTs), a new quantum workforce with special qualification will be needed. Building this workforce requires educational efforts, from short-term training to degree programs. In order to plan, map and compare such efforts, personal qualifications or job requirements, standardization is necessary. The European Competence Framework for Quantum Technologies (CFQT) provides a common language for QT education. The 2024 update to version 2.5 extends it with the new proficiency triangle and qualification profiles: The proficiency triangle proposes six proficiency levels for three proficiency areas, specifying knowledge and skills for each level. Nine qualification profiles show prototypical qualifications or job roles relevant to the quantum industry with the required proficiency, examples, and suggestions. This is an important step towards the standardization of QT education. The CFQT update is based on the results of an analysis of 34 interviews on industry needs. The initial findings from the interviews were complemented by iterative refinement and expert consultation, as well as by ML-driven analysis and mapping of job posts, revealing recommendations for formulating job posts.
keywords:
quantum technologies, competence framework, CFQT, qualification profiles, proficiency triangle, quantum industry, education, interview analysis, LLM
1 Introduction
Modern Quantum Technologies (QTs), such as quantum computers, quantum sensors or quantum communication devices, are gaining increasing industrial relevance. Building the future quantum workforce requires not only intensive educational efforts such as training and study programs. QT education also needs a common language as well as standardization to make efforts and qualifications comparable: The European Competence Framework for Quantum Technologies (CFQT) [1] is the “reference framework for planning, mapping and comparing QT-related educational activities, personal qualification and job requirements”. This paper documents the update from version 2.0 to 2.5 and the advances in the new version.
The CFQT aims to structure QT competences as well as related competences relevant to the future quantum workforce. It was already used, for example, to map course content [2], to categorize research [3], or as the backbone for a curriculum transformation framework [4]. Version 2.5 (2024) of the CFQT was published by the European Commission, EU Publications Office (see Ref. [1]). It was developed within the European Quantum Flagship [5] Coordination and Support Action (CSA) projects QTEdu [6] and QUCATS [7]. Within QUCATS, the European Quantum Readiness Center (EQRC) [8] has been launched to establish best practices, including standardization, using the CFQT as a tool.
Version 2.0 of the CFQT focused on content relevant in the context of QT, structured in the content map. The content map and the details for the eight content domains have remained consistent from version 2.0 (2023) to 2.5 (2024). With the update, two key additions have been made: the proficiency triangle (see Sec. 3.1) and the qualification profiles (see Sec. 3.2). Appendix A.1 provides an overview of the general structure of the CFQT document.
Crucial input for the CFQT extension in version 2.5 came from the analysis of 34 interviews focusing on QT educational needs in industry. The primary objective of the interview analysis was to report on the qualification and training needs in industry, as documented in Ref. [9]. However, the insights from the interviews provided great input for the CFQT update. In addition to the manual analysis, a GPT was employed to analyze anonymized interview transcripts and extract typical job roles to improve the initial qualification profiles. The GPT was also used to map job posts to the qualification profile descriptions and thus evaluate the improvement of the profile descriptions through three draft versions.
2 Methods and procedure for updating to version 2.5
The development of the CFQT started with an iterative study in 2020/2021 [10], leading to version 1.0 [11] compiled within the QTEdu CSA [6]. Within the QUCATS project, it was updated to version 2.0, as documented in Ref. [12]. The update to version 2.5, which brought the proficiency triangle and qualification profiles into the framework, was released in April 2024. The process and results of the update are documented below, including several excerpts from the CFQT document for illustration. A version history and related publications are given in Appendix A.2.
2.1 Interview analysis and iterative refinement
In summer 2023, 34 interviews with industry representatives were conducted and analyzed to answer research questions including What QT qualification and training needs are reported by industry? An in-depth analysis is documented in Ref. [9], including details about the methodology, which was based on qualitative content analysis [13]. The interviewees represent QT companies or companies with QT departments, from start-ups to very large companies, covering technical, management or business roles.
In some of the interviews, the material ‘competence types’ (Fig. 5) was shown to stimulate the flow of the interviews. This was an outline for structuring typical competences that we expected to be relevant in industry. Three levels were listed for each of the three ‘competence types’: build/develop QT (level B–Build…, C–Component improvement…, D–Develop…), overview/communicate about QT (level N, O, P), and use/adapt QT (U, V, W), see Appendix B. The key competences and qualifications discussed in the interviews were categorized according to this structure. However, it turned out that it was incomplete and not disjunctive, so refinement was needed.
Through an iterative process, this initial structure was refined into what we named the proficiency triangle, see Sec. 3.1. It consists of three proficiency areas with six proficiency levels each. For each of these 18 levels, a competence statement was formulated using action verbs [14]. They are complemented by descriptions of knowledge and skills whose formulations are based on the European Qualifications Framework (EQF) [15]. Already for the CFQT version 2.0 [16], the EQF was used to describe six proficiency levels, as documented in Ref. [12]. The proficiency level labels A1, A2, B1, B2, C1 and C2 are associated with the levels commonly used to specify qualifications in languages in Europe [17]. The DigCompEdu [18] framework was used as a template which also provided the initial version of the level keywords, e.g. awareness for level A1. Those were also updated to better match the level descriptions.
All proficiency levels were formulated in a generic way, adapted from the EQF definitions for QT in general. In the new version 2.5, they are specified for the three proficiency areas, and key competences identified in the interviews were incorporated.
In addition, categories related to job roles were formulated. During the qualitative content analysis of the interview transcripts, eleven categories were used, mainly based on the structure of the ‘competence types’ (Fig. 5). In Table 5, these job role categories are shown, together with a reference to the related ‘competence type’ level.
In the next step, 30 anonymized interview transcripts were analyzed using a large language model (LLM, namely GPT3.5) to identify frequently mentioned personas representing categories or clusters of job roles. This resulted in the descriptions of the 10 personas provided in Appendix C.1. These personas were checked against the initial job role categories, supporting the reduction from eleven to nine profile-related categories and an initial refinement. For the remaining categories, a second iteration of categorizing the interview passages was performed. Several categories in the second iteration are already named similarly to the final profiles, as documented in Table 5. Based on these categories, or more precisely, the categorized interview passages, the qualification profile descriptions version 2.1 were created.
These profiles were then refined through discussions with several colleagues and by gathering feedback from experts in both academia and industry. In this way, the titles and descriptions were refined, and example personas and suggestions were added based on the results of the interview analysis documented in Ref [9].
2.2 LLM-based evaluation of the qualification profiles
As a first trial of the consistency of the qualification profiles and the robustness of their descriptions, we used an LLM in order to test the profiles for stability under repeated classifications, adapting a methodology previously used for sentiment classification [19]. We used a dataset of 50 QT related job descriptions taken from a larger collection of LinkedIn job posts collected for research at the EQRC [8]. The job descriptions were classified into one of the nine qualification profiles using the LLM.
We ran the matching algorithm 15 times with the LLM gpt3.5-turbo-16k and calculated the stability of the classification as $\frac{x}{15}$, where $x$ is the most frequent classification into one profile for each job post, which is divided by the number of total classifications (15). For all 50 jobs, the mean value $s=\frac{1}{50}\sum\left(\frac{x}{15}\right)$ was calculated as a metric of the classification stability. For the three versions of the profile descriptions, these are $s_{1}=0.944$ (v2.1), $s_{2}=0.927$ (v2.2) and $s_{3}=0.981$ (v2.3). This demonstrates an increase from the first version to the third version. The full distribution of profile classifications is visualized in Fig. 6, data are available on Zenodo [20].
In all versions, the majority of the jobs were stable under 15 classifications, as measured by the stability $s$. However, some jobs were split across classes. For version 2.3, three jobs descriptions were less than 85% stable under repeated classifications (i.e., the most frequent classification $x$ was less than 13 within 15 repetitions, thus IDs 10, 37 and 40 in Fig. 6). In order to understand this, these job descriptions were examined individually. This has implications for how job posts should be written in order to most clearly communicate the role, which we consider in Sec. 3.3.
2.3 Limitations
The update of the CFQT is primarily based on the analysis of 34 interviews with participants from the quantum industry. This is a strength of the dataset, as it is directly constructed from interviews with experts in the field. However, we should anticipate and acknowledge that there may be a bias due to the group of interviewees, as discussed in Ref. [9]. Most of the perspectives included come from people who have a very positive perception of QT, both in the interviews and in the iterative refinement. Therefore, they may identify personas more readily than those without direct involvement in the industry would.
The extraction of personas from the interview transcripts by an LLM improves the objectivity of the qualitative content analysis by having a kind of second coder analyze most of the interviews (the 30 transcripts that could be anonymized) and prepare a second set of categories.
In addition, the QT field is very active and rapidly changing. As the interviews were conducted in 2023, the identified needs that influenced the CFQT update may also change in a short period of time. To capture these dynamic developments, annual updates of the CFQT are planned as part of the QUCATS project, with the next one due in April 2025, and feedback is taken continuously.
3 Results: Proficiency triangle, qualification profiles and implications from the job post mapping
3.1 Proficiency triangle: 3 proficiency areas with 6 levels each
The new proficiency triangle adds a second dimension, proficiency, to the content map from the CFQT version 2.0. It consists of three proficiency areas (I) quantum concepts, (II) QT hardware and software (HW & SW) engineering, and (III) QT applications & strategies. The proficiency areas are described in the CFQT document, for example:
Area (III) QT applications & strategies addresses the business dimension of QT applications (5.6, 6.7, 7.4, 7.6). It focuses on the question of “how to generate value with QT” (domain 8). This also includes considerations of impact and responsibility, extending to the exploration of novel applications and the design of new products utilizing QT. [quoted from 1, p. 14]
The proficiency of an individual may grow in these three areas independently, however, typically someone working in the QT industry would have proficiency in all three areas. For each of these three proficiency areas, six proficiency levels A1, A2, B1, B2, C1 and C2 are specified. Together they form the proficiency triangle shown in Fig. 1.
(I) Quantum concepts(II) QT HW & SW engineering(III) QT applications & strategies
Figure 1: The three proficiency areas (left) and their visualization in the proficiency triangle from the CFQT [1, p. 30] (right). The proficiency levels are labeled A1, A2, B1, B2, C1, and C2, with A levels representing beginner, B levels intermediate and C levels advanced. Proficiency first grows in breadth up to level A2, preparing for specialization in the B levels. At the tip of the triangle (level C2), proficiency reaches and extends the state of the art, with strong specialization. However, proficiency also grows continuously in breadth, so that additional, e.g. rather B1 level, proficiency contributes to the higher levels, e.g. level C2.
With the coverage of the triangle, a QT-specific set of knowledge and skills, or a qualification can be visualized. The qualification profiles show prototypical coverage of the proficiency triangle: Each profile is visualized by a partially colored proficiency triangle, indicating the level of proficiency with respect to each of the three proficiency areas, see Sec. 3.2.
Each proficiency level consists of a competence statement, as well as knowledge and skills descriptions based on the EQF level definitions (as described in Sec. 2.1). The competence statements are reproduced in Table 1.
Table 1: Competence statements (short version) from the CFQT [1, p. 14]. There are three additional pages in the CFQT that provide more detail on the proficiency levels. A ‘QT facet’ stands for a QT core, component, system, or application which can be HW and/or SW focused.
Proficiency level Area (I) Area (II) Area (III)
A1 Awareness Reproduce basic q. concepts & terminology Reproduce basic functionalities of a QT facet Recognize potential of QT
A2 Literacy Describe fundamental q. concepts Perform basic tasks on a QT facet Identify value of QT
B1 Utilization Apply quantum methods to problems Modify/apply a QT facet Classify available QT applications/approaches
B2 Investigation Analyze problems with quantum Analyze performance, improve QT Analyze QT market and opportunities
C1 Specialization Refine and extend quantum methods Conceptualize integrated QT systems Advise on QT appl. selection or strategies
C2 Innovation Develop innovative solutions Develop new QT facet Develop and assess QT (product) strategies
In addition, for each of the proficiency levels, relations to the content map (sub)domains (5.6, 6.7, …) and examples of how the levels could be reached are given. For example, the description of the lowest proficiency level A1 for the proficiency area (III) QT applications & strategies is:
A1 Awareness: Recognize potential of QT
  • K:
    Basic idea of the potential of QT systems and applications, overview of possibilities, challenges and limitations.
  • S:
    Ability to follow public media and discussions with critical awareness of hype.
with the related content (sub)domains & examples:
Applications for a QT, see 5.6, 6.7, 7.4, 7.6, or a selected (sub)topic (concrete application area, e.g. quantum optimization in logistics 5.6), with relevance for business 8.2 or education 8.4; hype (especially for computing).
Multiple choice questions on possibilities/limitations. [quoted from 1, p. 17]
Examples of how to test or measure whether someone has reached this level are included (in italics). They are a first step towards a certification scheme. The proficiency level descriptions are formulated independently of, e.g., a concrete QT. To specify, e.g., a specific qualification of a person or the aim of a course or the requirements for a job, the proficiency triangle must be combined with a selection from the content map. For each level, there are references to related content (sub)domains. In the example above these are:
  • 5.6
    applications of quantum computing and simulation, including material science, engineering and design, optimization …   [p. 10]
  • 6.7
    applications of quantum sensors, including metrology at single quantum level, medicine and molecular biology, …  [p. 11]
  • 7.4, 7.6
    applications of quantum cryptography …; system networks (composite systems), quantum internet applications, … [p. 12]
  • 8.2
    business strategy, entrepreneurship and management (business level), including risk management and strategic decision making, … [p. 13]
  • 8.4
    responsibility and awareness, including ethics, public communication and outreach, education … [p. 13]
All these subdomains of the content map influence the ‘flavor’ of an individual’s experience or the learning outcomes of a course, i.e. the specific technologies or concepts covered. A course may cover only applications in quantum computing (subdomain 5.6), or even only quantum optimization in logistics and thus only (sub)topics from subdomain 5.6 regarding applications. However, to cover level A1 in the business-related proficiency area (III), the relevance to business (8.2) needs to be covered with a critical perspective, e.g. addressing the hype in quantum computing. The focus would be on the value and impact of QT for an industry sector or the own company/business.
Alternatively, the focus could be on the value and impact for society and education, i.e. subdomain 8.4 instead of 8.2. With 8.4 one can map qualifications outside of industry, e.g. a school teacher who recognizes the potential of QT as a context for teaching quantum concepts and who has an overview of how QT can impact society as a whole. This would be relevant, e.g., to discuss QT and their expected impact already in high school and thus to get more school students interested in quantum (and STEM). Raising this interest was rated as very necessary (40 of 52 participants rated ‘high need’, 11 rated ‘low need’, one rated ‘no need’) in our follow-up survey to the interview analysis, see Ref. [9, Sec. 5.4].
3.2 Qualification profiles
The qualification of a person or by a training can be described by a coverage of the proficiency triangle, which specifies the proficiency level for each of the three proficiency areas. Since these are formulated independently of the concrete subject matter, they have to be combined with a selection from the content map. Nine qualification profiles included in the CFQT version 2.5 show the prototypical coverage of the proficiency triangle. They are listed in Table 2 with the required proficiency levels for each of the three proficiency areas.
Table 2: Qualification profiles and related proficiency levels in the CFQT version 2.5.
Proficiency level in area
  (I)  (II) (III)
No. Profile title q. backgr. QT eng. appl./strat.
P1 QT aware person A1 A1 A1
P2 QT informed decision maker A1
P3 QT literate person (QT literate business role, advocator, enthusiast) A2 A2 A2
P4 QT practitioner (working with QT, technician, QT user) A1 B1 A1
P5 QT business analyst A2 A2 B2
P6 QT engineering professional (e.g. QT engineer, quantum computer or information scientist) B1 B2 A2
P7 QT (HW or SW) specialist (e.g. senior QT engineer, QT architect) B2 C2 B1
P8 QT (product) strategist (e.g. advisor, business development expert) B2 B2 C2
P9 QT core innovator C2 B1 A2
An overview page shows these nine profiles with the different, partly colored proficiency triangles as well as the relations between the profiles [1, p. 18]. Relations mean how people may evolve from one profile to another. As in the example in Fig. 2, such a potential evolution is visualized by arrows.
Figure 2: Three qualification profiles with profile relations, from the CFQT summary [1, p. 30]. The arrows visualize the profile relation, the solid line shows a typical progression, the dashed line shows another conceivable qualification path. The color of the lines corresponds to the color of the proficiency area in which the qualification grows from one profile to another.
For each profile, a detail page provides the partially colored proficiency triangle with the descriptions of the highest proficiency level covered for each of the three proficiency areas. The profile is described generally, and example personas provide more concrete descriptions of what a job with that qualification might look like, for example for profile 4:
General description
The QT practitioner is someone who works around the development, assembly and operation of QT (technicians or ‘classical’ engineers with some QT specific additional qualifications), or uses QT with some customization:
  • is QT aware, thus is able to follow team discussions and has an idea of the potential of the QT working on,
  • has an overview of the relevant parts (hardware and/or software) for QT, and
  • focuses on the specific QT relevant to their own work, and knows how to work with it.
Example personas
‘Classical’ engineer with QT add-on, QT lab technician (e.g. for operation and maintenance), QT assembly and test technician, …
An engineer working on QT development, could be an electronic or mechanical engineer or a software engineer/computer scientist, working on the control hardware/software for a qubit, needs mainly traditional engineering skills, but works together with the quantum people, so needs an idea of the special challenges in QT development (but does not need to understand the details, has a supervisor who ensures compliance with quantum requirements) and also has an idea of the applications etc. to know what they are working for. […] [1, p. 22]
The profile description is complemented by a ‘needs and suggestions’ section, describing the suggested previous qualification (i.e. another profile and/or a specific background), training modules to reach the profile starting from the suggested previous qualification. In addition, recommendations regarding language and certificates are provided. All these suggestions are based on the interview analysis. Some of the results documented in Ref. [9] are included in these sections of the qualification profile detail pages. They include what training is considered relevant or what language is deemed appropriate.
Finally it should be emphasized that the qualification profiles – like the proficiency level description – are formulated independent of a specific technology in which they are applied. In order to describe a specific qualification, the qualification profile or, more generally, the coverage of the proficiency triangle must be combined with a topical selection from the content map. Two examples are given in the CFQT: the quantum control electronics practitioner for NV sensors and the quantum optimization in logistics & production analyst [1, p. 28 & 29]. Figure 3 shows the coverage of the content map along with the focus topics for the first of these examples. The clear focus on quantum sensing and control technologies is visible. They complement the partially colored proficiency triangle for the QT practitioner (P4, right triangle in Fig. 2): it shows the coverage of proficiency level A1 for the proficiency areas (I) and (III) along with level B1 in area (II). Each of the two examples is completed by a description of the example and the associated QT specific qualification.
Figure 3: Content map coverage and focus topics for the quantum control electronics practitioner for NV sensors [1, p. 28]. On the left side, the covered domains from the content map are depicted: domains 1, 3, 4, 6 and 8. While the domains 1 and 8 are only covered at a low level, as indicated in the proficiency triangle (right triangle for P4 in Fig. 2) with level A1 for proficiency area (I) and (III), the specialization is in area (II) up to level B1. This specialization is also shown on the right side (focus contents). Here, the selected subdomains with selected topics provide information about the concrete topics in the focus of the qualification.
3.3 Additional findings and implications from the job mapping
As discussed in Sec. 2.2, 50 job posts were classified by an LLM to three versions of the nine qualification profile descriptions. Most job posts were stable under repeated classification, visible in the statistical distribution shown in the Appendix, Fig. 6. Table 3 lists the titles of the qualification profiles in the three versions used for the job post mapping and the published version.
Table 3: Evolution of titles of the qualification profiles. Although the titles may not have changed between two iterations, the descriptions may have changed significantly.
No. v2.1 v2.2 v2.3 v2.5 (published in Ref [1])
1 QT aware person
2 QT aware decision maker QT informed decision maker
3 QT-literate communicator QT communicator QT literate person
4 QT (market) analyst QT business analyst renumbered to P5
5 QT user, engineer working with QT QT technician P4 QT practitioner
6 QT engineer (generalist) QT engineer/scientist (generalist) QT engineering professional
7 senior QT engineer/QT architect QT (HW or SW) specialist
8 QT strategist QT (product) strategist
9 QT core innovator
We notice that the jobs matching to the qualification profiles are not evenly distributed. For example, there are more job offers matching to profiles 5 (renumbered to profile 4 in the published version, see Table 3), 6, and 7 – the more technical and engineering roles – than those matching to the others. This is consistent with research currently being conducted within the QUCATS project at the EQRC [8, subpage analysis], which indicates that the majority of jobs in the QT field are currently technical, such as researchers or engineers, rather than non-technical, such as in business or management.
Taking a closer look on unstable job posts, we identified that these contained descriptions that did not represent the role as accurately as those with more stable classifications, or did not represent them at all, rather describing the company. For example, these unstable job descriptions included a) a software developer (ID 40): a very long (760 words) description of a company with no reference to the tasks of the role itself; b) an algorithm developer (ID 37), consisting of a description of the role of an engineering staff (profile 6) but also emphasizing tasks for a more senior position (profile 7); c) an open call for applicants of any role (ID 10, mapping to profiles v2.3).
As a practical result of the job post mapping, we conclude that not all job descriptions are completely suited to be mapped using the nine qualification profiles. We recommend that job descriptions be written with at least three specific sections: (a) Skills required to be qualified to perform the job, (b) tasks/responsibilities of the role, and (c) formal qualifications required to apply. Out of 1113 QT job posts aggregated by the EQRC [8] since January 2023, 895 (80.4%) are written with all three of these recommended sections. Therefore, roughly one in five job posts in the field is written with insufficient information to reliably profile the role. In order for graduates to be able to prepare for the job market, with full knowledge of the skills required for the position they are applying for, we recommend that employers consider including complete information in their job posts, enabling the role to be clearly understood.
4 Discussion and conclusion
Prior to implementing new courses and programs to develop the workforce pipeline in QT, it is critical to answer a key question: What specific knowledge and skills are needed for industrx growth? Different studies were conducted to analyze these needs [21, 22, 23, 24, 25, 10], including our analysis of 34 interviews with industry representatives [9] This research provided the primary input for the extension of the CFQT with the proficiency triangle and qualification profiles. The proficiency levels organized in the proficiency triangle cover key knowledge and skills discussed in the interviews, and the qualification profiles are supplemented by example personas and needs and suggestions extracted from the interview analysis.
The CFQT provides a common language for QT-related qualification and education. It is a tool for planning educational activities such as workshops, training or study programs. It can be used to specify the qualification of an individual, or what is required for a job. For example, it is useful for …
  • educators to plan educational activities and map their objectives;
  • individuals to specify what qualification they have and what they want to achieve;
  • learners to identify what training or course they should take;
  • companies to specify what qualifications an employee or team covers and compare that to what they would need to identify what qualifications are missing;
  • companies to specify job requirements and compare them with a candidate’s qualifications;
  • job seekers to map their qualifications and compare them to job requirements.
Early users are educators [e.g. 2, 4]. Two EU-funded projects agreed to use the framework to map their activities: DigiQ [26], developing 16 QT master’s programs across Europe, and QTIndu [27], developing industrial training courses on QT.
The CFQT is an important step towards the standardization of QT education. This need was also addressed in one of the interviews analyzed for the update to version 2.5:
I think it is always better to have a standard. So the question is how to standardize these trainings? […] It comes back to the ECTS systems from education. I mean, they are standards.  [quote from an interview with a QT start-up; 9, p. 24]
For the update to version 2.0, we already aligned the proficiency level descriptions with the EQF and thus the European Credit Transfer and Accumulation System (ECTS), as described in Ref. [12, p. 3]. Following the EQF [15, p. 16], the three highest EQF levels are associated with the first, second and third cycles of the Qualifications Framework of the European Higher Education Area, which are based on the ECTS. Therefore, as the CFQT levels are formulated based on the EQF. The level B2 is linked to bachelor, C1 to master and C2 to doctoral programs. However, it should be emphasized that this mapping is not necessarily one-to-one, and that a degree program is not the only way to reach these levels. Especially in the business-related area (III), the high levels are more likely to be achieved through years of experience.
In version 2.0 of the CFQT, these links were explicitly stated in the proficiency level descriptions. For example, it was noted for level B2: “e.g. through a short research project as for a bachelor thesis, internship with project” [16, p. 5]. As mentioned above, this brings the risk of assuming a study program as the only way to reach a level, while there is a variety of opportunities to gain such a qualification.
In version 2.5, such comments are less prominent and vary across the three proficiency areas. For the same proficiency level B2 Investigation, these are for area (I) “Short research project with documentation, e.g. student research project or bachelor thesis”, for area (II) “Write report on requirements analysis and performance results” and for area (III) “Document an analysis for a concrete potential use case covering potential advances, risks, […]” [1, p. 15-17]. With the update, the focus in these remarks shifts from an approximate duration to what an individual should be able to answer, perform or prepare for a qualification covering the corresponding proficiency level. Suggestions for training formats, the need for study programs or work experience are provided for the qualification profiles, rather than being fixed for a proficiency level.
The qualification profiles, together with the suggestions listed on the detail pages, are the starting point for a certification scheme to ensure comparability of industry training across Europe. This scheme is currently being developed within the Quantum Flagship coordination project, QUCATS [7]. As a part of QUCATS, the EQRC also highlights “accords” (best practices) of organizations across the EU [8, subpage Accords]. Within these accords, the CFQT is a tool that can provide structure and standardization. As a recent example, the accord submitted by the QuantUM Group at the University of Minho provides a profile of a software engineer in quantum computing as a representative graduate from the new master degree established there [8, subpage Accords: Educational Accords]. In addition, the EQRC website hosts a curated playlist of videos, sorted based on the content map of the CFQT [8, subpage Resources]. The CFQT, consisting of the content map and the proficiency triangle with associated qualification profiles, provides a reference framework for greater standardization across Europe. However, adoption is critical to its success. Industry representatives seeking new employees should consider using the CFQT to understand their needs. They should apply the CFQT to define skills, requirements and responsibilities in their job advertisements. Educational institutions also have an important role to play. Here, the qualification profiles can be a step towards the future of standardized learning outcomes in QTs.
Abbreviations
  • CFQT
    European Competence Framework for Quantum Technologies [1]
  • CSA
    Coordination and Support Action
  • DigiQ
    Digitally enhanced quantum technology master (project title) [26]
  • EQF
    European Qualifications Framework [15]
  • EQRC
    European Quantum Readiness Center [8]
  • GPT
    Generative Pre-trained Transformer (one type of LLMs)
  • HW
    Hardware
  • LLM
    Large Language Model
  • ML
    Machine Learning
  • QT/QTs
    Quantum Technology/-ies
  • QTEdu
    Quantum Technology Education (QTEdu CSA, project title) [6]
  • QTIndu
    Quantum Technologies courses for Industry (project title) [27]
  • QUCATS
    Quantum Flagship Coordination AcTion and Support (project title) [5]
  • STEM
    Science, Technology, Engineering, and Mathematics
  • SW
    Software
Declarations
[…]
Data availability The three versions v2.1-v2.3 of profile descriptions, the 50 job posts, the prompt and the mapping statistics are available on Zenodo, see Ref. [20]. Here, also the data from the industry needs analysis (Ref. [9]) is available.
[…]
Acknowledgements We thank Carolyn Levy, Dennis Green, Dion Timmermann, and all experts from industry and academia for their valuable feedback within the iterative refinement of the descriptions of the proficiency level or qualification profiles. FG prepared, conducted and analyzed the interviews, prepared the CFQT drafts, collected feedback and wrote the original manuscript. SG conducted the machine learning based analysis and wrote the related part of the original manuscript (Sec. 2.2 and 3.3). All authors contributed to the study design and analysis, i.e. they discussed the methods and results, revised the manuscript, and read and approved the final manuscript.
[…]
Funding This work has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101070193.
This publication reflects only the views of the authors, the European Commission is not responsible for any use that may be made of the information it contains.
[…]
Competing interests The authors have no competing interests to declare that are relevant to the content of this article.
[…]
Financial Interests The authors have no relevant financial or non-financial interests to disclose.
[…]
Ethics approval Ethical review and approval was not required for the study on human participants in accordance with the local legislation and institutional requirements.
Appendix A Structure and version history
A.1 Principal structure of the CFQT
The CFQT [1] consists of three main parts:
  1. 1.
    The content map provides a structured overview of QT related topics and contents. It is a kind of extended table of contents with up to four layers: domains 1, 2, …, 8 with 42 subdomains 1.1, 1.2, …, (p. 5 in the CFQT document) and, on the detail pages for each domain (p. 6–13), additional topics and sometimes subtopics provide more details for the subdomains. A selection of domains from the content map and of subdomains with topics are included in an example shown in Fig. 3.
  2. 2.
    The proficiency triangle visualizes six proficiency levels for three proficiency areas (p. 14; Fig. 1). In-depth level descriptions and more are provided for each of the three proficiency areas (p. 15-17). See Sec. 3.1.
  3. 3.
    The qualification profiles show typical qualification, i.e. coverage of the proficiency triangle. An overview (p. 18) and detailed descriptions including example personas and (training) suggestions (p. 19–27) are available. See Sec. 3.2.
Furthermore, two examples how to combine a profile with a content selection are given (p. 28–29 in the CFQT). In addition, the CFQT pages 3 and 4 provide some information on how to use the framework, including an introduction of the terminology and the level system, and information on the methodology and related publications. The coloring of the proficiency triangle corresponds to the coloring of the content map as shown in Fig. 4.
Figure 4: Coloring relation between content map (left) and proficiency triangle (right) [1, p. 4].
A.2 Version history and related publications
For the Qualification Profiles, a beta version was published in 2022 [28] as a separate document based on the CFQT version 1.0. The new profiles replace this beta version, they are not an update of the beta version and were created with a completely different approach. While four of the six profiles in the beta version were those of an engineer in QT hardware or one of the QT pillars, the new profiles focus on the proficiency that a QT engineer needs in general. These four beta version profiles would be the same profile in the new version. Now the pillar is the context to be added by a topical selection of the content map. In the beta version, the profiles consisted mainly of the topical selection of the content map, and only provided a very limited specification with proficiency levels, which themselves were not specified at that time.
Version 2.5 of the CFQT was released in April 2024 on Zenodo [29] and in August 2024 as an ‘official version’ by the Publications Office of the European Union [1]. They differ only in the placement or addition of some editorial information (mainly moved from page 3 to the new page 1), but the framework itself has not changed between the EU and Zenodo publications. Table 4 provides a version history and an overview about related publications.
Table 4: Version history and publications related to the CFQT, with CFQT standing for ‘European Competence Framework for Quantum Technologies’ and ‘QT(s)’ for ‘Quantum Technology(-is)’ (in the titles). [*] Older versions of the CFQT are also available on Zenodo [29].
Date [Ref.] Title/version Comment
Dec. 2020 [30] Beta version of the CFQT Content map with seven domains and details pages, one keyword for each of six proficiency levels. Available on request on Zenodo.
June 2022
(Nov. 2020)
[31] Requirements for future quantum workforce – a Delphi study Interim report on the iterative study (Ref. [10]), conference proceedings of GIREP, Nov. 2020.
May 2021 CFQT version 1.0 (before graphical update) Update of content map, restructuring the QT pillar domains to a total of eight domains.
Sep. 2021 [*] CFQT version 1.0 Only graphical update.
Sep. 2021 [11] Competence framework for quantum technologies: methodology and version history Documentation of the development process until version 1.0, published by the EU Publications Office.
Jan. 2022 [28] Qualification Profiles for QTs (beta version) Six profiles: content map selection with proficiency level indication, outdated.
June 2023
(Aug. 2022)
[10] Future quantum workforce: Competences, requirements, and forecasts Paper on the iterative study collecting input for the CFQT beta version and also version 1.0, submitted in Aug. 2022
April 2023 [*] CFQT version 2.0 Update of content map, addition of descriptions for six proficiency levels.
July 2023 [12] Towards a quantum ready workforce: the updated CFQT Paper on the update to v2.0.
April 2024 [29] CFQT version 2.5 (Zenodo) Extension by proficiency triangle, new qualification profiles, content map unchanged.
July 2024 [9] Advancing QT workforce: industry insights into qualification and training needs Preprint on the analysis of interviews with industry that were also used to update the CFQT to version 2.5
Aug. 2024 [1] CFQT version 2.5 Official published by EU publications office.
Appendix B From competence types through categories to qualification profiles
In some of the interviews (22 out of 34), the material ‘Competence types’ was shown, reproduced in Fig. 5. They formed the initial system that was refined and extended through the CFQT update process, as described in Sec. 2.1.
Figure 5: Competence types as a material shown in the interviews, June 2023, outdated. For better readability, the descriptions are reproduced below.
The three competence types shown in Fig. 5 with three levels each are:
Overview, communicate (overview on applications and use cases)
  • N
    Notice the basic idea and potential of an application and (possible) use cases.
  • O
    Overview of an application’s landscape and use cases, critical perspective on potential and limitations, know and communicate with basic quantum vocabulary.
  • P
    Perform a comparison of different applications based on a deeper understanding of the various functions, assess which application is suitable for which use case.
Build, develop (build application)
  • B
    Build an application component, assemble.
  • C
    Component improvement for an application.
  • D
    Develop new components or applications through research, selection and integration of components.
Use, adapt (use an application)
  • U
    Use/run an application for one specific use case that is customized for that use case.
  • V
    Variate/adapt an application for several related use cases.
  • W
    Wrap applications and components together to enable new use cases.
As described in Ref. [9], the 34 interviews with industry representatives were analyzed using qualitative content analysis. Table 5 shows an extract from the category system related to the qualification profiles (left column) as well as a revised version of these categories used for a second iteration of the categorization (middle column). In addition, the qualification profiles in the CFQT version 2.5 are listed (right column). For the first iteration, also the related ‘competence type’ level is given, e.g. [B] for ‘Build an application…’, as specified above. The progression from competence type levels to qualification profiles is visible in rows.
Table 5: Category system and statistic for the two iterations related to the development of the qualification profiles together with the final profile number (No.) title in the CFQT version 2.5. In the columns called N, the number of segments categorized in the category is given. If no number is given, the code is a supportive code only to structure the codes on the next level.
First iteration Second iteration Qualification Profile
Category N Category N No. Title in CFQT
quantum aware workforce
 basic user (click a button) [U] 6 other: not quantum specific/ no quantum needed 16
 engineers with (almost) no quantum [B] 22 QT aware person 18 P1 QT aware person
 hype, basic idea (not) for all, incl. admin people [N] 24 QT aware decision maker, manager 17 P2 QT informed decision maker
quantum literate workforce
 management (non-QT comp.), business, sales, policy makers [N/O] 42 QT literate (sales) communicator 27 P3 QT literate person
 use-case identifier [O/W] 19  use case identification 5
 eng. with overview on QT (also e.g. cryo physicists; incl. programming) [O/B] 73 QT user, engineer working with QT 69 P4 QT practitioner
 advanced user (with adaption) [V] 23  QT user (with adaptation) 19
q. expert workforce (needs QT study program or PhD)
 strategist, consultants [P] (comparison, assessment) 27 QT analyst (market, use cases, strategies) 15 P5 QT business analyst
QT engineer (generalist) 6 P6 QT engineering professional
senior QT engineer, QT architect P7 QT (HW or SW) specialist
 customer/sales eng. [C/V] 12  … focus on customer and applications (incl. consulting, perform comparison) 21
 quantum algorithm developers (high-level) [D] 22   QT application specialist (e.g. programming) 22
 systems eng. (integration, industrialization) [C] 13  … focus on systems and integration 12
QT strategist (e.g. consultant) 2 P8 QT (product) strategist
 quantum experts/
expertise needs [D]
35 QT core innovator 29 P9 QT core innovator
Appendix C From LLM: personas and stability testing
C.1 Extracted personas
As described in Sec. 2.1, 30 anonymized interview transcripts were provided to an LLM to extract ten personas as typical job roles. These are:
  1. 1.
    Quantum Engineers/Researchers/Scientists: These individuals are involved in the design and development of quantum technologies, requiring expertise in quantum physics, engineering, and material science.
  2. 2.
    Project Managers: Responsible for overseeing the successful development and implementation of quantum-enabled devices, coordinating efforts between different teams or departments.
  3. 3.
    Decision Makers/Management: Individuals in managerial or executive roles responsible for making strategic decisions based on market needs and technology trends in the quantum industry.
  4. 4.
    Training and Development Professionals: Facilitate the learning process for existing employees or new hires, designing hybrid training programs, online content, workshops, or lectures.
  5. 5.
    Generalists: Individuals with a broader understanding of quantum technologies, bridging the gap between technical and market-oriented aspects of quantum technologies.
  6. 6.
    Market Analysts: Assess market demand and guide the development of quantum-enabled devices by understanding market trends, potential applications, and customer needs.
  7. 7.
    Company Representatives: Engaged in technology development, consulting, and strategic planning for quantum technology companies, addressing workforce development and skilled personnel needs.
  8. 8.
    End-users: Users in various industries, such as medical imaging, interested in understanding and benefiting from quantum-enhanced technologies.
  9. 9.
    Technical Developers: Involved in building physical devices and components for quantum technologies, such as lasers and high-precision optical and frequency components.
  10. 10.
    Sales Managers and Marketing Personnel: Responsible for understanding the market landscape, guiding sales engineers, shaping product strategy, and communicating the value of quantum technologies to potential customers.
These personas were used to refine the initial categories from the qualitative content analysis of the interviews (see Table 5) and thus influenced the preparation of version 2.1 of the qualification profile descriptions. The profile descriptions were iteratively refined. This refinement was accompanied by a mapping of job descriptions: 50 job descriptions were categorized into one of the nine profile descriptions for versions 2.1, 2.2, and 2.3, and a stability test (see Fig. 6) was performed.
C.2 Stability testing
Figure 6 shows the stability testing for the mapping of 50 job posts. Table 3 contains the titles of the profiles in the three versions used for the stability testing, the full descriptions (version 2.1, 2.2, 2.3) are available on Zenodo, Ref [20], along with the job descriptions used for the stability testing, the prompt and the classification statistics.
Figure 6: Stability testing of 50 jobs (row ID) under 15 classifications into one of nine qualification profiles, using QP description versions 2.1, 2.2 and 2.3, and GPT model gpt3.5-turbo-16k.

### v3

@@
 With the increasing industrial relevance of quantum technologies (QTs), a new quantum workforce with special qualification will be needed.
-Building this workforce requires educational efforts, from short-term training to degree programs.
+Building this workforce requires educational efforts, ranging from short term training to degree programs.
 In order to plan, map and compare such efforts, personal qualifications or job requirements, standardization is necessary.
 The European Competence Framework for Quantum Technologies (CFQT) provides a common language for QT education.
-The 2024 update to version 2.5 extends it with the new proficiency triangle and qualification profiles: The proficiency triangle proposes six proficiency levels for three proficiency areas, specifying knowledge and skills for each level.
-Nine qualification profiles show prototypical qualifications or job roles relevant to the quantum industry with the required proficiency, examples, and suggestions.
+The 2024 update to version 2.5 extends it by the new proficiency triangle and qualification profiles: The proficiency triangle proposes six proficiency levels for three proficiency areas, specifying knowledge and skills for each level.
+Nine qualification profiles show prototypical qualifications or job roles relevant to the quantum industry, with the required proficiency, examples, and suggestions.
 This is an important step towards the standardization of QT education.
 The CFQT update is based on the results of an analysis of 34 interviews on industry needs.
-The initial findings from the interviews were complemented by iterative refinement and expert consultation, as well as by ML-driven analysis and mapping of job posts, revealing recommendations for formulating job posts.
+The initial findings from the interviews were complemented by iterative refinement and expert consultation.
 keywords:
@@
 Building the future quantum workforce requires not only intensive educational efforts such as training and study programs.
-QT education also needs a common language as well as standardization to make efforts and qualifications comparable: The European Competence Framework for Quantum Technologies (CFQT) [1] is the “reference framework for planning, mapping and comparing QT-related educational activities, personal qualification and job requirements”.
-This paper documents the update from version 2.0 to 2.5 and the advances in the new version.
+QT education also needs a common language as well as standardization efforts to make education and qualifications comparable: The European Competence Framework for Quantum Technologies (CFQT) [1] is the “reference framework for planning, mapping and comparing QT-related educational activities, personal qualification and job requirements”.
+This paper documents the update from version 2.0 to version 2.5 and the advances in the new version.
 The CFQT aims to structure QT competences as well as related competences relevant to the future quantum workforce.
-It was already used, for example, to map course content [2], to categorize research [3], or as the backbone for a curriculum transformation framework [4].
+It was already used, for example, to map course content [2], to plan new courses [3], to categorize research [4], or as the backbone for a curriculum transformation framework [5].
 Version 2.5 (2024) of the CFQT was published by the European Commission, EU Publications Office (see Ref.
@@
 Within QUCATS, the European Quantum Readiness Center (EQRC) [9] has been launched to establish best practices, including standardization, using the CFQT as a tool.
-Version 2.0 of the CFQT focused on content relevant in the context of QT, structured in the content map.
-The content map and the details for the eight content domains have remained consistent from version 2.0 (2023) to 2.5 (2024).
-With the update, two key additions have been made: the proficiency triangle (see Sec.
-3.1) and the qualification profiles (see Sec.
-3.2).
-Appendix A.1 provides an overview of the general structure of the CFQT document.
+The objective of the update to version 2.5 was to provide more detailed and substantiated descriptions for the proficiency levels, as well as a new version for the qualification profiles from 2022 [beta version, 10].
+General descriptions for six proficiency levels were already added in the CFQT update to version 2.0, but needed further refinement and specification for QT – as in the new proficiency triangle (see Sec.
+3).
+The proficiency triangle enables the specification of proficiency or qualification in three dimensions, covering (I) quantum concepts, thus physics fundamentals, (II) QT engineering competences, including QT functionalities, and (III) QT applications and strategies, i.e.
+the business perspective, including aspects such as impact and ethics.
+By assigning a proficiency level for each of these three proficiency areas, the qualification of an individual can be specified, for example.
+A total of nine new qualification profiles were identified and visualized with the proficiency triangle (see Sec.
+4).
+They represent prototypical qualifications relevant to industry.
+However, to really specify such a qualification, a second measure is needed: A selection from the content map.
+It structures content relevant in the context of QT and was in the focus of the update to version 2.0 [11].
+The content map and the details for the eight content domains remained consistent from version 2.0 (2023) to 2.5 (2024).
+In the new version, two examples are given of how to combine such a selection from the content map and a profile based on the proficiency triangle.
+Appendix A.1 provides an overview of the general structure of the CFQT document, from the content map and the proficiency triangle to the qualification profiles and examples.
 Crucial input for the CFQT extension in version 2.5 came from the analysis of 34 interviews focusing on QT educational needs in industry.
@@
 [12].
-However, the insights from the interviews provided great input for the CFQT update.
+However, the insights from the interviews provided great input for the CFQT update: key competences were identified and incorporated into the proficiency level descriptions, and the jobs and roles discussed in the interviews led to the first draft of the new qualification profiles.
 In addition to the manual analysis, a GPT was employed to analyze anonymized interview transcripts and extract typical job roles to improve the initial qualification profiles.
-The GPT was also used to map job posts to the qualification profile descriptions and thus evaluate the improvement of the profile descriptions through three draft versions.
 2 Methods and procedure for updating to version 2.5
 The development of the CFQT started with an iterative study in 2020/2021 [13], leading to version 1.0 [14] compiled within the QTEdu CSA [7].
-Within the QUCATS project, it was updated to version 2.0, as documented in Ref.
+Within the QUCATS project [8], it was updated to version 2.0, as documented in Ref.
 [11].
+In this update, the content map was revised and descriptions of six proficiency levels A1 to C2 were added.
+The proficiency level descriptions were based on the European Qualifications Framework (EQF) [15], thus specifying the required knowledge and skills for each proficiency level adapted to QT in general.
+The proficiency level labels A1, A2, B1, B2, C1 and C2 are associated with the levels commonly used to specify qualifications in languages in Europe [16].
+The DigCompEdu [17] framework was used as a template which also provided the initial version of the level keywords, e.g.
+awareness for level A1.
+These keywords were also updated in version 2.5 to better match the level descriptions.
 The update to version 2.5, which brought the proficiency triangle and qualification profiles into the framework, was released in April 2024.
-The process and results of the update are documented below, including several excerpts from the CFQT document for illustration.
+The process and results of the update are documented below, including some excerpts from the CFQT document for illustration.
 A version history and related publications are given in Appendix A.2.
-2.1 Interview analysis and iterative refinement
+2.1 About the interviews: introduction of ‘competence types’
 In summer 2023, 34 interviews with industry representatives were conducted and analyzed to answer research questions including What QT qualification and training needs are reported by industry?
@@
 [12], including details about the methodology, which was based on qualitative content analysis [18].
-The interviewees represent QT companies or companies with QT departments, from start-ups to very large companies, covering technical, management or business roles.
-In some of the interviews, the material ‘competence types’ (Fig.
-5) was shown to stimulate the flow of the interviews.
+The interviewees represent QT companies or companies with QT departments from different European countries, from start-ups to very large companies, covering technical, management or business roles.
+In some of the interviews, a slide about the ‘competence types’ (reproduced in Fig.
+7) was shown to stimulate the flow of the interviews – others were conducted without this input, e.g.
+if the interviewee expressed very clear or different views in the previous questions.
+This material introduced a structure of competence descriptions for three levels for each of the three ‘types’ overview, communicate (level N, O, P), build, develop (level B, C, D) and use, adapt (level U, V, W):
+Overview, communicate (overview on applications and use cases)
+  • N
+    Notice the basic idea and potential of an application and (possible) use cases.
+  • O
+    Overview of an application’s landscape and use cases, critical perspective on potential and limitations, know and communicate with basic quantum vocabulary.
+  • P
+    Perform a comparison of different applications based on a deeper understanding of the various functions, assess which application is suitable for which use case.
+Build, develop (build application)
+  • B
+    Build an application component, assemble.
+  • C
+    Component improvement for an application.
+  • D
+    Develop new components or applications through research, selection and integration of components.
+Use, adapt (use an application)
+  • U
+    Use/run an application for one specific use case that is customized for that use case.
+  • V
+    Variate/adapt an application for several related use cases.
+  • W
+    Wrap applications and components together to enable new use cases.
 This was an outline for structuring typical competences that we expected to be relevant in industry.
-Three levels were listed for each of the three ‘competence types’: build/develop QT (level B–Build…, C–Component improvement…, D–Develop…), overview/communicate about QT (level N, O, P), and use/adapt QT (U, V, W), see Appendix B.
-The key competences and qualifications discussed in the interviews were categorized according to this structure.
-However, it turned out that it was incomplete and not disjunctive, so refinement was needed.
-Through an iterative process, this initial structure was refined into what we named the proficiency triangle, see Sec.
-3.1.
+Thus, the key competences and qualifications discussed in the interviews were categorized according to this structure, regardless of whether the material was used in the interview or not.
+It turned out that this structure was incomplete and not disjunctive, so refinement was needed.
+More concretely, the structure of the competence types was the starting point for developing or extracting both, the proficiency triangle and the qualification profiles, as discussed in the next sections and visualized in Fig.
+1: The proficiency triangle addresses the issue of disjunctive proficiency areas with independent proficiency levels for each area.
+In contrast, the qualification profiles show typical combinations of these levels, visualized as coverage of the proficiency triangle, for a more complete picture than the ‘competence types’ provided.
+Figure 1: From the ‘competence types’ (top) to the proficiency triangle (left) and the qualification profiles (right), which use the proficiency triangle as visualization.
+The issues of the ‘competence types’ of being incomplete and not disjunctive descriptions of competences are addressed through the proficiency triangle.
+The second issue that rather combinations of competency type levels were discussed in industry is addressed through the qualification profiles.
+2.2 From ‘competence types’ to the proficiency triangle
+Key competences were identified in the interview transcripts and analyzed not only in terms of the corresponding proficiency level, but also in terms of the three categories, which were initially the three ‘competence types’ that iteratively evolved into the three proficiency areas:
+  • from “overview […] communicate with basic quantum vocabulary”, plus additions to area (I) quantum concepts,
+  • from ‘build, develop’ and ‘use, adapt’ to (II) QT hardware (HW) & software (SW) engineering, and
+  • aspects from ‘overview, communicate’, e.g. “critical perspective on potential and limitations”, plus additions to (III) QT applications & strategies.
+Through another iterative process, the graphical appearance of a triangle was constructed.
+In this way, the initial structure of the ‘competence types’ was refined into what we named the proficiency triangle, see Sec.
+3.
 It consists of three proficiency areas with six proficiency levels each.
-For each of these 18 levels, a competence statement was formulated using action verbs [14].
-They are complemented by descriptions of knowledge and skills whose formulations are based on the European Qualifications Framework (EQF) [15].
-Already for the CFQT version 2.0 [16], the EQF was used to describe six proficiency levels, as documented in Ref.
-[12].
-The proficiency level labels A1, A2, B1, B2, C1 and C2 are associated with the levels commonly used to specify qualifications in languages in Europe [17].
-The DigCompEdu [18] framework was used as a template which also provided the initial version of the level keywords, e.g.
-awareness for level A1.
-Those were also updated to better match the level descriptions.
-All proficiency levels were formulated in a generic way, adapted from the EQF definitions for QT in general.
-In the new version 2.5, they are specified for the three proficiency areas, and key competences identified in the interviews were incorporated.
-In addition, categories related to job roles were formulated.
-During the qualitative content analysis of the interview transcripts, eleven categories were used, mainly based on the structure of the ‘competence types’ (Fig.
-5).
-In Table 5, these job role categories are shown, together with a reference to the related ‘competence type’ level.
-In the next step, 30 anonymized interview transcripts were analyzed using a large language model (LLM, namely GPT3.5) to identify frequently mentioned personas representing categories or clusters of job roles.
-This resulted in the descriptions of the 10 personas provided in Appendix C.1.
-These personas were checked against the initial job role categories, supporting the reduction from eleven to nine profile-related categories and an initial refinement.
-For the remaining categories, a second iteration of categorizing the interview passages was performed.
+For each proficiency level, a competence statement was formulated using action verbs [19].
+They are complemented by descriptions of knowledge and skills based on the EQF levels [15], as were the general level descriptions in version 2.0 [20, 11].
+In contrast to the descriptions in version 2.0, the proficiency levels are now specified for the three proficiency areas and are based on the key competences identified through the interview analysis.
+Further details, such as relations to the content map, are also provided.
+All details for one proficiency level are given as an example in Sec.
+3.
+2.3 From the identification of job roles to qualification profiles
+During the qualitative content analysis of the interview transcripts, twelve categories related to job roles were used.
+These categories were mainly based on the structure of the ‘competence types’ (Sec.
+2.1/Fig.
+7), as well as some frequently discussed combinations of two ‘competence type’ levels as additional categories.
+In Appendix B.2, these job role categories are shown, together with references to related ‘competence type’ levels.
+For example, the category “engineer with overview on QT [O/B]” is related to the “O – Overview…” and “B – Build…” and evolved to profile P4 of the QT practitioner, including also the “advanced user [V]”, related to “V – Variate…”.
+Iteratively, the categories from the interview analysis evolved to the qualification profiles.
+To support this process, 30 anonymized interview transcripts were analyzed using a large language model (LLM, namely GPT3.5) to identify frequently mentioned job roles representing categories or clusters.
+This resulted in the descriptions of the 10 roles provided in Appendix B.3.
+These roles were checked against the initial job role categories, in addition to intensive discussion of the initial categories.
+On this basis, the job role categories were refined, and a second iteration of categorizing the interview passages was performed.
 Several categories in the second iteration are already named similarly to the final profiles, as documented in Table 5.
@@
 These profiles were then refined through discussions with several colleagues and by gathering feedback from experts in both academia and industry.
-In this way, the titles and descriptions were refined, and example personas and suggestions were added based on the results of the interview analysis documented in Ref [9].
-2.2 LLM-based evaluation of the qualification profiles
-As a first trial of the consistency of the qualification profiles and the robustness of their descriptions, we used an LLM in order to test the profiles for stability under repeated classifications, adapting a methodology previously used for sentiment classification [19].
-We used a dataset of 50 QT related job descriptions taken from a larger collection of LinkedIn job posts collected for research at the EQRC [8].
-The job descriptions were classified into one of the nine qualification profiles using the LLM.
-We ran the matching algorithm 15 times with the LLM gpt3.5-turbo-16k and calculated the stability of the classification as $\frac{x}{15}$, where $x$ is the most frequent classification into one profile for each job post, which is divided by the number of total classifications (15).
-For all 50 jobs, the mean value $s=\frac{1}{50}\sum\left(\frac{x}{15}\right)$ was calculated as a metric of the classification stability.
-For the three versions of the profile descriptions, these are $s_{1}=0.944$ (v2.1), $s_{2}=0.927$ (v2.2) and $s_{3}=0.981$ (v2.3).
-This demonstrates an increase from the first version to the third version.
-The full distribution of profile classifications is visualized in Fig.
-6, data are available on Zenodo [20].
-In all versions, the majority of the jobs were stable under 15 classifications, as measured by the stability $s$.
-However, some jobs were split across classes.
-For version 2.3, three jobs descriptions were less than 85% stable under repeated classifications (i.e., the most frequent classification $x$ was less than 13 within 15 repetitions, thus IDs 10, 37 and 40 in Fig.
-6).
-In order to understand this, these job descriptions were examined individually.
-This has implications for how job posts should be written in order to most clearly communicate the role, which we consider in Sec.
-3.3.
+For example, the profile names were changed to better represent industry roles, and the required proficiency levels for the profiles were discussed and adjusted based on these discussions.
+In addition to invited discussions with individual experts, an open event was conducted in March 2024 to gather feedback from the QTEdu community [7, subpage Events].
+Table 1 shows the evolution of the profile titles from version 2.1 to the published version.
+The titles and descriptions were refined, and example personas/job roles and suggestions were added based on the results of the interview analysis documented in Ref [12].
+Table 1: Evolution of titles of the qualification profiles.
+Although the titles may not have changed between two iterations, the descriptions may have changed significantly.
+These descriptions of the profiles in versions 2.1, 2.2, and 2.3 are available in the supplementary material on Zenodo [21].
+No. v2.1 v2.2 v2.3 v2.5 (published in Ref [1])
+1 QT aware person
+2 QT aware decision maker QT informed decision maker
+3 QT-literate communicator QT communicator QT literate person
+4 QT (market) analyst QT business analyst renumbered to Profile 5
+5 QT user, engineer working with QT QT technician Profile 4 QT practitioner
+6 QT engineer (generalist) QT engineer/scientist (generalist) QT engineering professional
+7 senior QT engineer/QT architect QT (HW or SW) specialist
+8 QT strategist QT (product) strategist
+9 QT core innovator
 2.4 Limitations
@@
 Most of the perspectives included come from people who have a very positive perception of QT, both in the interviews and in the iterative refinement.
-Therefore, they may identify personas more readily than those without direct involvement in the industry would.
-The extraction of personas from the interview transcripts by an LLM improves the objectivity of the qualitative content analysis by having a kind of second coder analyze most of the interviews (the 30 transcripts that could be anonymized) and prepare a second set of categories.
+Therefore, they may identify QT-related job roles more readily than those without direct involvement in the industry would.
+The extraction of job roles from the interview transcripts by an LLM improves the objectivity of the qualitative content analysis by having a kind of second coder analyze most of the interviews (the 30 transcripts that could be anonymized) and prepare a second set of categories.
+However, the experts involved in the refinement of the qualification profiles must be expected to have a similar bias as the interviewees, since they have a similar background and experience.
 In addition, the QT field is very active and rapidly changing.
@@
 To capture these dynamic developments, annual updates of the CFQT are planned as part of the QUCATS project, with the next one due in April 2025, and feedback is taken continuously.
-3 Results: Proficiency triangle, qualification profiles and implications from the job post mapping
-3.1 Proficiency triangle: 3 proficiency areas with 6 levels each
+3 Proficiency triangle
 The new proficiency triangle adds a second dimension, proficiency, to the content map from the CFQT version 2.0.
-It consists of three proficiency areas (I) quantum concepts, (II) QT hardware and software (HW & SW) engineering, and (III) QT applications & strategies.
+It consists of three proficiency areas (I) quantum concepts, (II) QT HW & SW engineering, and (III) QT applications & strategies.
 The proficiency areas are described in the CFQT document, for example:
@@
 Together they form the proficiency triangle shown in Fig.
-1.
+2.
+With the coverage of the triangle, a QT-specific set of knowledge and skills, or a qualification can be visualized, as in the qualification profiles, see Sec.
+4.
 (I) Quantum concepts(II) QT HW & SW engineering(III) QT applications & strategies
@@
 level C2.
-With the coverage of the triangle, a QT-specific set of knowledge and skills, or a qualification can be visualized.
-The qualification profiles show prototypical coverage of the proficiency triangle: Each profile is visualized by a partially colored proficiency triangle, indicating the level of proficiency with respect to each of the three proficiency areas, see Sec.
-3.2.
 Each proficiency level consists of a competence statement, as well as knowledge and skills descriptions based on the EQF level definitions (as described in Sec.
-2.1).
+2.2).
 The competence statements are reproduced in Table 2.
@@
 Proficiency level Area (I) Area (II) Area (III)
-A1 Awareness Reproduce basic q. concepts & terminology Reproduce basic functionalities of a QT facet Recognize potential of QT
-A2 Literacy Describe fundamental q. concepts Perform basic tasks on a QT facet Identify value of QT
+A1 Awareness Reproduce basic quantum concepts & terminology Reproduce basic functionalities of a QT facet Recognize potential of QT
+A2 Literacy Describe fundamental quantum concepts Perform basic tasks on a QT facet Identify value of QT
 B1 Utilization Apply quantum methods to problems Modify/apply a QT facet Classify available QT applications/approaches
@@
 In addition, for each of the proficiency levels, relations to the content map (sub)domains (5.6, 6.7, …) and examples of how the levels could be reached are given.
-For example, the description of the lowest proficiency level A1 for the proficiency area (III) QT applications & strategies is:
+For example, the description of the lowest proficiency level A1 for the proficiency area (III) QT applications & strategies is [quoted from 1, p. 17]:
 A1 Awareness: Recognize potential of QT
@@
     Ability to follow public media and discussions with critical awareness of hype.
-with the related content (sub)domains & examples:
+For each level, related content (sub)domains & examples are given, in this case:
 Applications for a QT, see 5.6, 6.7, 7.4, 7.6, or a selected (sub)topic (concrete application area, e.g.
 quantum optimization in logistics 5.6), with relevance for business 8.2 or education 8.4; hype (especially for computing).
-Multiple choice questions on possibilities/limitations. [quoted from 1, p. 17]
-Examples of how to test or measure whether someone has reached this level are included (in italics).
-They are a first step towards a certification scheme.
-The proficiency level descriptions are formulated independently of, e.g., a concrete QT.
+Examples of how to test or measure whether someone has reached a proficiency level are also provided.
+For this example, it is: Multiple choice questions on possibilities/limitations.
+These remarks are a first step toward a certification scheme that will specify the measurement of proficiency in more detail.
+The proficiency level descriptions are formulated independently of, e.g., a concrete QT pillar.
 To specify, e.g., a specific qualification of a person or the aim of a course or the requirements for a job, the proficiency triangle must be combined with a selection from the content map.
 For each level, there are references to related content (sub)domains.
-In the example above these are:
-  • 5.6
-    applications of quantum computing and simulation, including material science, engineering and design, optimization …   [p. 10]
-  • 6.7
-    applications of quantum sensors, including metrology at single quantum level, medicine and molecular biology, …  [p. 11]
-  • 7.4, 7.6
-    applications of quantum cryptography …; system networks (composite systems), quantum internet applications, … [p. 12]
-  • 8.2
-    business strategy, entrepreneurship and management (business level), including risk management and strategic decision making, … [p. 13]
-  • 8.4
-    responsibility and awareness, including ethics, public communication and outreach, education … [p. 13]
+In the example above, these are the subdomains 5.6, 6.7, 7.4, 7.6, 8.2 and 8.4.
+They are visualized in Fig.
+3.
+Figure 3: Related subdomains for proficiency level A1 in proficiency area (III), extracted from the content map [1, p. 5].
 All these subdomains of the content map influence the ‘flavor’ of an individual’s experience or the learning outcomes of a course, i.e.
@@
 Raising this interest was rated as very necessary (40 of 52 participants rated ‘high need’, 11 rated ‘low need’, one rated ‘no need’) in our follow-up survey to the interview analysis, see Ref.
-[9, Sec.
-5.4].
-3.2 Qualification profiles
-The qualification of a person or by a training can be described by a coverage of the proficiency triangle, which specifies the proficiency level for each of the three proficiency areas.
+[12, p.
+25].
+4 Qualification profiles
+The qualification of an individual or the objectives of a training, for example, can be visualized by a coverage of the proficiency triangle, which specifies the proficiency level for each of the three proficiency areas.
 Since these are formulated independently of the concrete subject matter, they have to be combined with a selection from the content map.
@@
 They are listed in Table 3 with the required proficiency levels for each of the three proficiency areas.
-Table 2: Qualification profiles and related proficiency levels in the CFQT version 2.5.
+Table 3: Qualification profiles and related proficiency levels for area (I) quantum concepts, (II) QT HW & SW engineering and (III) QT applications and strategies in the CFQT version 2.5.
 Proficiency level in area
-  (I)  (II) (III)
-No. Profile title q. backgr. QT eng. appl./strat.
+No. Profile title   (I)  (II) (III)
 P1 QT aware person A1 A1 A1
 P2 QT informed decision maker A1
-P3 QT literate person (QT literate business role, advocator, enthusiast) A2 A2 A2
+P3 QT literate person (QT lit. business role, advocator, enthusiast) A2 A2 A2
 P4 QT practitioner (working with QT, technician, QT user) A1 B1 A1
@@
 As in the example in Fig.
-2, such a potential evolution is visualized by arrows.
+4, such a potential evolution is visualized by arrows.
 Figure 4: Three qualification profiles with profile relations, from the CFQT summary [1, p.
@@
 For each profile, a detail page provides the partially colored proficiency triangle with the descriptions of the highest proficiency level covered for each of the three proficiency areas.
-The profile is described generally, and example personas provide more concrete descriptions of what a job with that qualification might look like, for example for profile 4:
-General description
-The QT practitioner is someone who works around the development, assembly and operation of QT (technicians or ‘classical’ engineers with some QT specific additional qualifications), or uses QT with some customization:
+The profile is described generally, and example personas/job roles provide more concrete descriptions of what a job with that qualification might look like, for example for profile 4:
+General description The QT practitioner is someone who works around the development, assembly and operation of QT (technicians or ‘classical’ engineers with some QT specific additional qualifications), or uses QT with some customization:
   • is QT aware, thus is able to follow team discussions and has an idea of the potential of the QT working on,
@@
   • focuses on the specific QT relevant to their own work, and knows how to work with it.
-Example personas
-‘Classical’ engineer with QT add-on, QT lab technician (e.g. for operation and maintenance), QT assembly and test technician, …
-An engineer working on QT development, could be an electronic or mechanical engineer or a software engineer/computer scientist, working on the control hardware/software for a qubit, needs mainly traditional engineering skills, but works together with the quantum people, so needs an idea of the special challenges in QT development (but does not need to understand the details, has a supervisor who ensures compliance with quantum requirements) and also has an idea of the applications etc.
+Example personas (‘Classical’ engineer with QT add-on, QT lab technician (e.g. for operation and maintenance), QT assembly and test technician, …)
+An
+engineer working on QT development, could be an electronic or mechanical engineer or a software engineer/computer scientist, working on the control hardware/software for a qubit, needs mainly traditional engineering skills, but works together with the quantum people, so needs an idea of the special challenges in QT development (but does not need to understand the details, has a supervisor who ensures compliance with quantum requirements) and also has an idea of the applications etc.
 to know what they are working for.
@@
 They complement the partially colored proficiency triangle for the QT practitioner (P4, right triangle in Fig.
-2): it shows the coverage of proficiency level A1 for the proficiency areas (I) and (III) along with level B1 in area (II).
+4): it shows the coverage of proficiency level A1 for the proficiency areas (I) and (III) along with level B1 in area (II).
 Each of the two examples is completed by a description of the example and the associated QT specific qualification.
@@
 While the domains 1 and 8 are only covered at a low level, as indicated in the proficiency triangle (right triangle for P4 in Fig.
-2) with level A1 for proficiency area (I) and (III), the specialization is in area (II) up to level B1.
+4) with level A1 for proficiency area (I) and (III), the specialization is in area (II) up to level B1.
 This specialization is also shown on the right side (focus contents).
 Here, the selected subdomains with selected topics provide information about the concrete topics in the focus of the qualification.
-3.3 Additional findings and implications from the job mapping
-As discussed in Sec.
-2.2, 50 job posts were classified by an LLM to three versions of the nine qualification profile descriptions.
-Most job posts were stable under repeated classification, visible in the statistical distribution shown in the Appendix, Fig.
-6.
-Table 3 lists the titles of the qualification profiles in the three versions used for the job post mapping and the published version.
-Table 3: Evolution of titles of the qualification profiles. Although the titles may not have changed between two iterations, the descriptions may have changed significantly.
-No. v2.1 v2.2 v2.3 v2.5 (published in Ref [1])
-1 QT aware person
-2 QT aware decision maker QT informed decision maker
-3 QT-literate communicator QT communicator QT literate person
-4 QT (market) analyst QT business analyst renumbered to P5
-5 QT user, engineer working with QT QT technician P4 QT practitioner
-6 QT engineer (generalist) QT engineer/scientist (generalist) QT engineering professional
-7 senior QT engineer/QT architect QT (HW or SW) specialist
-8 QT strategist QT (product) strategist
-9 QT core innovator
-We notice that the jobs matching to the qualification profiles are not evenly distributed.
-For example, there are more job offers matching to profiles 5 (renumbered to profile 4 in the published version, see Table 3), 6, and 7 – the more technical and engineering roles – than those matching to the others.
-This is consistent with research currently being conducted within the QUCATS project at the EQRC [8, subpage analysis], which indicates that the majority of jobs in the QT field are currently technical, such as researchers or engineers, rather than non-technical, such as in business or management.
-Taking a closer look on unstable job posts, we identified that these contained descriptions that did not represent the role as accurately as those with more stable classifications, or did not represent them at all, rather describing the company.
-For example, these unstable job descriptions included a) a software developer (ID 40): a very long (760 words) description of a company with no reference to the tasks of the role itself; b) an algorithm developer (ID 37), consisting of a description of the role of an engineering staff (profile 6) but also emphasizing tasks for a more senior position (profile 7); c) an open call for applicants of any role (ID 10, mapping to profiles v2.3).
-As a practical result of the job post mapping, we conclude that not all job descriptions are completely suited to be mapped using the nine qualification profiles.
-We recommend that job descriptions be written with at least three specific sections: (a) Skills required to be qualified to perform the job, (b) tasks/responsibilities of the role, and (c) formal qualifications required to apply.
-Out of 1113 QT job posts aggregated by the EQRC [8] since January 2023, 895 (80.4%) are written with all three of these recommended sections.
-Therefore, roughly one in five job posts in the field is written with insufficient information to reliably profile the role.
-In order for graduates to be able to prepare for the job market, with full knowledge of the skills required for the position they are applying for, we recommend that employers consider including complete information in their job posts, enabling the role to be clearly understood.
 5 Discussion and conclusion
-Prior to implementing new courses and programs to develop the workforce pipeline in QT, it is critical to answer a key question: What specific knowledge and skills are needed for industrx growth?
-Different studies were conducted to analyze these needs [21, 22, 23, 24, 25, 10], including our analysis of 34 interviews with industry representatives [9] This research provided the primary input for the extension of the CFQT with the proficiency triangle and qualification profiles.
-The proficiency levels organized in the proficiency triangle cover key knowledge and skills discussed in the interviews, and the qualification profiles are supplemented by example personas and needs and suggestions extracted from the interview analysis.
+Prior to implementing new courses and programs to develop the workforce pipeline in QT, it is critical to answer a key question: What specific knowledge and skills are needed for industry growth?
+Different studies were conducted to analyze these needs [22, 23, 24, 25, 26, 13], including our analysis of 34 interviews with industry representatives [12].
+This research provided the primary input for the extension of the CFQT with the proficiency triangle and qualification profiles.
+The proficiency levels organized in the proficiency triangle cover key knowledge and skills discussed in the interviews, and the qualification profiles are supplemented by example personas/job roles and needs and suggestions extracted from the interview analysis.
 The CFQT provides a common language for QT-related qualification and education.
@@
 Early users are educators [e.g.
-2, 4].
+2, 5, 3].
 Two EU-funded projects agreed to use the framework to map their activities: DigiQ [27], developing 16 QT master’s programs across Europe, and QTIndu [28], developing industrial training courses on QT.
+Additionally, first companies showed interest in using the CFQT for their workforce development.
 The CFQT is an important step towards the standardization of QT education. This need was also addressed in one of the interviews analyzed for the update to version 2.5:
@@
 I mean, they are standards.
-[quote from an interview with a QT start-up; 9, p.
+[quote from an interview with a QT start-up; 12, p.
 24]
 For the update to version 2.0, we already aligned the proficiency level descriptions with the EQF and thus the European Credit Transfer and Accumulation System (ECTS), as described in Ref.
-[12, p.
+[11, p.
 3].
@@
 16], the three highest EQF levels are associated with the first, second and third cycles of the Qualifications Framework of the European Higher Education Area, which are based on the ECTS.
-Therefore, as the CFQT levels are formulated based on the EQF.
-The level B2 is linked to bachelor, C1 to master and C2 to doctoral programs.
+Therefore, as the CFQT levels are formulated based on the EQF, the level B2 is linked to bachelor, C1 to master and C2 to doctoral programs.
 However, it should be emphasized that this mapping is not necessarily one-to-one, and that a degree program is not the only way to reach these levels.
-Especially in the business-related area (III), the high levels are more likely to be achieved through years of experience.
+Especially in the business-related proficiency area (III), the high levels are more likely to be achieved through years of experience.
 In version 2.0 of the CFQT, these links were explicitly stated in the proficiency level descriptions.
 For example, it was noted for level B2: “e.g.
-through a short research project as for a bachelor thesis, internship with project” [16, p.
+through a short research project as for a bachelor thesis, internship with project” [20, p.
 5].
@@
 15-17].
-With the update, the focus in these remarks shifts from an approximate duration to what an individual should be able to answer, perform or prepare for a qualification covering the corresponding proficiency level.
+With the update, the focus in these remarks shifted from an approximate duration to what an individual should be able to answer, perform or prepare for a qualification covering the corresponding proficiency level.
 Suggestions for training formats, the need for study programs or work experience are provided for the qualification profiles, rather than being fixed for a proficiency level.
-The qualification profiles, together with the suggestions listed on the detail pages, are the starting point for a certification scheme to ensure comparability of industry training across Europe.
-This scheme is currently being developed within the Quantum Flagship coordination project, QUCATS [7].
-As a part of QUCATS, the EQRC also highlights “accords” (best practices) of organizations across the EU [8, subpage Accords].
-Within these accords, the CFQT is a tool that can provide structure and standardization.
-As a recent example, the accord submitted by the QuantUM Group at the University of Minho provides a profile of a software engineer in quantum computing as a representative graduate from the new master degree established there [8, subpage Accords: Educational Accords].
-In addition, the EQRC website hosts a curated playlist of videos, sorted based on the content map of the CFQT [8, subpage Resources].
+The updated CFQT is the starting point for a certification scheme to ensure comparability of industry training across Europe.
+This scheme is currently being developed within the Quantum Flagship coordination project, QUCATS [8], associated to standardization activities.
+As in technology standardization, the first step is typically to address terminology, already at very low technology readiness, followed by test and measurement standards [29].
+Similarly, with the CFQT we started to focus on the common language for QT education, based on the structured topics in the content map as well as the proficiency level descriptions.
+In version 2.5, there are already short, additional remarks on the aspect of testing or measuring if someone has reached a certain level.
+This aspect will be further elaborated in the certification scheme, a second step towards the standardization of QT education and qualification.
+As a part of QUCATS, the EQRC highlights “accords” (best practices) of organizations across the EU [9, subpage Accords].
+Within these accords, the CFQT is a tool to provide structure and standardization.
+As a recent example, the accord submitted by the QuantUM Group at the University of Minho provides a profile of a software engineer in quantum computing as a representative graduate from the new master degree established there [9, subpage Accords: Educational Accords].
+In addition, the EQRC website hosts a curated playlist of videos, sorted based on the content map of the CFQT [9, subpage Resources].
 The CFQT, consisting of the content map and the proficiency triangle with associated qualification profiles, provides a reference framework for greater standardization across Europe.
 However, adoption is critical to its success.
-Industry representatives seeking new employees should consider using the CFQT to understand their needs.
-They should apply the CFQT to define skills, requirements and responsibilities in their job advertisements.
+Industry representatives seeking new employees should consider using the CFQT to understand their needs and to formulate the required QT qualifications in their job advertisements.
 Educational institutions also have an important role to play.
 Here, the qualification profiles can be a step towards the future of standardized learning outcomes in QTs.
-Abbreviations
-  • CFQT
-    European Competence Framework for Quantum Technologies [1]
-  • CSA
-    Coordination and Support Action
-  • DigiQ
-    Digitally enhanced quantum technology master (project title) [26]
-  • EQF
-    European Qualifications Framework [15]
-  • EQRC
-    European Quantum Readiness Center [8]
-  • GPT
-    Generative Pre-trained Transformer (one type of LLMs)
-  • HW
-    Hardware
-  • LLM
-    Large Language Model
-  • ML
-    Machine Learning
-  • QT/QTs
-    Quantum Technology/-ies
-  • QTEdu
-    Quantum Technology Education (QTEdu CSA, project title) [6]
-  • QTIndu
-    Quantum Technologies courses for Industry (project title) [27]
-  • QUCATS
-    Quantum Flagship Coordination AcTion and Support (project title) [5]
-  • STEM
-    Science, Technology, Engineering, and Mathematics
-  • SW
-    Software
-Declarations
-[…]
-Data availability The three versions v2.1-v2.3 of profile descriptions, the 50 job posts, the prompt and the mapping statistics are available on Zenodo, see Ref.
-[20].
-Here, also the data from the industry needs analysis (Ref.
-[9]) is available.
-[…]
-Acknowledgements We thank Carolyn Levy, Dennis Green, Dion Timmermann, and all experts from industry and academia for their valuable feedback within the iterative refinement of the descriptions of the proficiency level or qualification profiles.
-FG prepared, conducted and analyzed the interviews, prepared the CFQT drafts, collected feedback and wrote the original manuscript.
-SG conducted the machine learning based analysis and wrote the related part of the original manuscript (Sec.
-2.2 and 3.3).
-All authors contributed to the study design and analysis, i.e.
-they discussed the methods and results, revised the manuscript, and read and approved the final manuscript.
-[…]
-Funding This work has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101070193.
-This publication reflects only the views of the authors, the European Commission is not responsible for any use that may be made of the information it contains.
-[…]
-Competing interests The authors have no competing interests to declare that are relevant to the content of this article.
-[…]
-Financial Interests The authors have no relevant financial or non-financial interests to disclose.
-[…]
-Ethics approval Ethical review and approval was not required for the study on human participants in accordance with the local legislation and institutional requirements.
+Further research is needed on the application of the CFQT.
+Whether it is used to prepare educational activities or within workforce development, accompanying research will be needed to identify issues and approaches for improving the CFQT.
+In addition, regular updates will be needed to reflect technological changes and innovations as well as new challenges and needs in the quantum industry.
 Appendix A Structure and version history
@@
 A selection of domains from the content map and of subdomains with topics are included in an example shown in Fig.
-3.
+5.
   2. 2.
@@
 14; Fig.
-1).
+2).
 In-depth level descriptions and more are provided for each of the three proficiency areas (p.
@@
 See Sec.
-3.1.
+3.
   3. 3.
@@
 See Sec.
-3.2.
+4.
 Furthermore, two examples how to combine a profile with a content selection are given (p.
@@
 The coloring of the proficiency triangle corresponds to the coloring of the content map as shown in Fig.
-4.
+6.
 Figure 6: Coloring relation between content map (left) and proficiency triangle (right) [1, p. 4].
@@
 April 2023 [*] CFQT version 2.0 Update of content map, addition of descriptions for six proficiency levels.
-July 2023 [12] Towards a quantum ready workforce: the updated CFQT Paper on the update to v2.0.
+July 2023 [11] Towards a quantum ready workforce: the updated CFQT Paper on the update to version 2.0.
 April 2024 [30] CFQT version 2.5 (Zenodo) Extension by proficiency triangle, new qualification profiles, content map unchanged.
@@
 Appendix B From competence types through categories to qualification profiles
-In some of the interviews (22 out of 34), the material ‘Competence types’ was shown, reproduced in Fig.
-5.
-They formed the initial system that was refined and extended through the CFQT update process, as described in Sec.
-2.1.
-Figure 5: Competence types as a material shown in the interviews, June 2023, outdated. For better readability, the descriptions are reproduced below.
-The three competence types shown in Fig. 5 with three levels each are:
-Overview, communicate (overview on applications and use cases)
-  • N
-    Notice the basic idea and potential of an application and (possible) use cases.
-  • O
-    Overview of an application’s landscape and use cases, critical perspective on potential and limitations, know and communicate with basic quantum vocabulary.
-  • P
-    Perform a comparison of different applications based on a deeper understanding of the various functions, assess which application is suitable for which use case.
-Build, develop (build application)
-  • B
-    Build an application component, assemble.
-  • C
-    Component improvement for an application.
-  • D
-    Develop new components or applications through research, selection and integration of components.
-Use, adapt (use an application)
-  • U
-    Use/run an application for one specific use case that is customized for that use case.
-  • V
-    Variate/adapt an application for several related use cases.
-  • W
-    Wrap applications and components together to enable new use cases.
+B.1 Competence types
+In some of the interviews (22 out of 34), the slide/material ‘competence types’ was shown, reproduced in Fig.
+7.
+They formed the initial system that was refined and extended to the proficiency triangle through the CFQT update process.
+Figure 7: Competence types as a material shown in the interviews, June 2023, outdated. For better readability, the descriptions are reproduced in Sec. 2.1.
+B.2 Categories and qualification profiles
 As described in Ref.
-[9], the 34 interviews with industry representatives were analyzed using qualitative content analysis.
+[12], the interviews 34 interviews with industry representatives were analyzed using qualitative content analysis.
 Table 5 shows an extract from the category system related to the qualification profiles (left column) as well as a revised version of these categories used for a second iteration of the categorization (middle column).
@@
 For the first iteration, also the related ‘competence type’ level is given, e.g.
-[B] for ‘Build an application…’, as specified above.
+[B] for ‘Build an application…’, as specified in Fig.
+7/Sec.
+2.1.
 The progression from competence type levels to qualification profiles is visible in rows.
-Table 5: Category system and statistic for the two iterations related to the development of the qualification profiles together with the final profile number (No.) title in the CFQT version 2.5.
+Table 5: Category system and statistics for the two iterations related to the development of the qualification profiles together with the final profile number (No.) title in the CFQT version 2.5.
 In the columns called N, the number of segments categorized in the category is given.
@@
  basic user (click a button) [U] 6 other: not quantum specific/ no quantum needed 16
- engineers with (almost) no quantum [B] 22 QT aware person 18 P1 QT aware person
- hype, basic idea (not) for all, incl. admin people [N] 24 QT aware decision maker, manager 17 P2 QT informed decision maker
+ engineers with (almost) no quantum [B] 22
+ hype, basic idea (not) for all, incl. admin people [N] 24 QT aware person 18 P1 QT aware person
+QT aware decision maker, manager 17 P2 QT informed decision maker
 quantum literate workforce
@@
 35 QT core innovator 29 P9 QT core innovator
-Appendix C From LLM: personas and stability testing
-C.1 Extracted personas
-As described in Sec. 2.1, 30 anonymized interview transcripts were provided to an LLM to extract ten personas as typical job roles. These are:
+B.3 GPT extracted job roles
+As described in Sec. 2.3, 30 anonymized interview transcripts were provided to an LLM (gpt3.5-turbo-16k) to extract ten personas as typical job roles. The prompts were:
+Prompt 1:
+Here is an anonymised interview transcript. What are the key personas described in the interview? (e.g Engineers, Developers, Project Managers)?
+The outputs of prompt 1, for each transcript, were collated into a single list.
+Prompt 2:
+This document contains the key personas described in a series of interviews with companies in the Quantum Technology industry.
+Each interview begins with the text “INAXXX”.
+Based on these findings, please provide a list of 10 types of personas in the Quantum Technology industry, each with a 2 sentence description.
+The extracted roles are:
   1. 1.
@@
 These personas were used to refine the initial categories from the qualitative content analysis of the interviews (see Table 5) and thus influenced the preparation of version 2.1 of the qualification profile descriptions.
-The profile descriptions were iteratively refined.
-This refinement was accompanied by a mapping of job descriptions: 50 job descriptions were categorized into one of the nine profile descriptions for versions 2.1, 2.2, and 2.3, and a stability test (see Fig.
-6) was performed.
-C.2 Stability testing
-Figure 6 shows the stability testing for the mapping of 50 job posts.
-Table 3 contains the titles of the profiles in the three versions used for the stability testing, the full descriptions (version 2.1, 2.2, 2.3) are available on Zenodo, Ref [20], along with the job descriptions used for the stability testing, the prompt and the classification statistics.
-Figure 6: Stability testing of 50 jobs (row ID) under 15 classifications into one of nine qualification profiles, using QP description versions 2.1, 2.2 and 2.3, and GPT model gpt3.5-turbo-16k.
+Abbreviations
+  • CFQT
+    European Competence Framework for Quantum Technologies [1]
+  • CSA
+    Coordination and Support Action
+  • DigiQ
+    Digitally enhanced quantum technology master (project title) [27]
+  • EQF
+    European Qualifications Framework [15]
+  • EQRC
+    European Quantum Readiness Center [9]
+  • HW/SW
+    Hardware/Software
+  • LLM
+    Large Language Model, e.g., GPT (Generative Pre-trained Transformer)
+  • QT/QTs
+    Quantum Technology/-ies
+  • QTEdu
+    Quantum Technology Education (QTEdu CSA, project title) [7]
+  • QTIndu
+    Quantum Technologies courses for Industry (project title) [28]
+  • QUCATS
+    Quantum Flagship Coordination AcTion and Support (project title) [6]
+  • STEM
+    Science, Technology, Engineering, and Mathematics
+Declarations
+[…]
+Data availability The three versions v2.1-v2.3 of profile descriptions are available on Zenodo [21]. Here, also data from the industry needs analysis (Ref. [12]) is available.
+[…]
+Acknowledgements We thank Carolyn Levy, Dennis Green, Dion Timmermann, and all experts from industry and academia for their valuable feedback within the iterative refinement of the descriptions of the proficiency level or qualification profiles.
+[…]
+Author contributions FG prepared, conducted and analyzed the interviews, prepared the CFQT drafts, collected feedback and wrote the original manuscript.
+SG conducted the LLM-based analysis (Sec.
+2.3).
+All authors contributed to the study design and analysis, i.e.
+they discussed the methods and results, revised the manuscript, and read and approved the final manuscript.
+[…]
+Funding This work has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101070193.
+This publication reflects only the views of the authors, the European Commission is not responsible for any use that may be made of the information it contains.
+[…]
+Competing interests The authors have no competing interests to declare that are relevant to the content of this article.
+[…]
+Financial Interests The authors have no relevant financial or non-financial interests to disclose.
+[…]
+Ethics approval Ethical review and approval was not required for the study on human participants in accordance with the local legislation and institutional requirements.